OpenStudy (anonymous):

Find the Absolute Value with solution set. then Graph 1.) 2x-3y=2 4x+5y=26 2.) x+y-6=0 3x-5y-34=0

terenzreignz (terenzreignz):

Absolute Value with solution set? I'm afraid I don't understand that :D

OpenStudy (anonymous):

maybe its my writing, but can you help me find the absolute value of both? and then graph

terenzreignz (terenzreignz):

Absolute value? Not sure what that is... as for the solution, I guess I can help out there...

terenzreignz (terenzreignz):

Quite frankly, I don't know what these mean... you're using an entirely different method...

terenzreignz (terenzreignz):

Not one that I've ever heard of :D

OpenStudy (anonymous):

5x-3y-16=0 4x-5y-5=0 Solution Set x| 2 |5 5(2)-36-16 5(5)-3y-16=0 y|-2 |3 10-3y-16 25-36-16=0 3y=16-10 36=16-25 3y=6 3y=9 3 3 y=-2 3 3 y=3 x|5|0 4(5)-5y-5=0 -5y-5=0 y|3|-1 20-5y=0 -5y=5 -5y=5-20 -5 -5 y=-1 -5=-15 -5 -5 y=3

OpenStudy (anonymous):

can you help me get the absolute value of this 5x-3y-16=0 4x-5y-5=0?

terenzreignz (terenzreignz):

Absolute value? I don't know what that means in this context.

OpenStudy (anonymous):

i cant seem to understand how they get this x| 2 |5 & x|5|0 y|-2 |3 y|3|-1

terenzreignz (terenzreignz):

LOL me neither... this is a method I am unfamiliar with... I guess the best I could do is attempt to solve it in the conventional way?

OpenStudy (anonymous):

and then graph?

terenzreignz (terenzreignz):

Yeah, graphing lines takes practice, but it's easy to get used to.

OpenStudy (anonymous):

ok then

terenzreignz (terenzreignz):

Well, can you solve the system? Using substitution.

terenzreignz (terenzreignz):

Faulty... though I admire your effort... extraordinary :D

OpenStudy (anonymous):

lol i think, i really need to find the absolute value 1st, thats what my teacher recommend to use.

terenzreignz (terenzreignz):

You'll need to consult with someone else for that.. I really don't know that method... sorry T.T

OpenStudy (anonymous):

ok then, switching to google :D

OpenStudy (anonymous):

but can you show me how to graph this?

terenzreignz (terenzreignz):

Well, graphing lines involves finding two points on that line, and then just connecting them to form a line :)

terenzreignz (terenzreignz):

It can actually be any two points, but the simplest is finding what are called the "intercepts". Heard that word before?

OpenStudy (anonymous):

nope

terenzreignz (terenzreignz):

Okay... an x-intercept is the point where the line intersects the x-axis. a y-intercept is the point where the lint intersects the y-axis. To get the x-intercept simply set y = 0, and solve for x. The x-intercept would be (x , 0) but you have to solve for x.... To get the y-intercept, set x = 0, and solve for y. The y-intercept would be (0 , y), after you solve for y.